### v2

@@ 4.3.1 Ablation Studies of ProCNS
Table 2: Ablation study on the three key components in ProCNS regarding DSC. Here, $\checkmark$ indicates the corresponding component is applied. “***” represents $\rm p\leq 0.05$ in a Wilcoxon matched-pairs signed rank test comparing the DSC of the pertinent component before and after application. The column for ODOC reports the average DSC of OD and OC. The best results are in bold.
@@
$\checkmark$ $*{}^{*}$ 87.91 $*{}^{*}$
$\checkmark$ $\checkmark$ $*{}^{*}$ 88.05 80.55
$\checkmark$ $\checkmark$ $\checkmark$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
To ascertain the efficacy of the three key components $\mathcal{L}_{\rm PRSA}$, ANPM and $\mathcal{L}_{\rm noise}$ within ProCNS, as well as their respective contributions to the overall pipeline, we perform a series of ablation studies. In this context, the baseline is defined as the segmentation model that is trained with sparse annotations and solely employs the pCE loss.Then, we incrementally incorporate the three components into the baseline. Table 2 presents the quantitative results from the ablation analyses on the three tasks in terms of DSC. Compared to the baseline, $\mathcal{L}_{\rm PRSA}$ respectively improves model performance by 2.36%, 0.52% and 0.47% for the FAZ, ODOC and polyp segmentation tasks. The integration of ANPM yields DSC improvements of 0.92%, 0.14% and 1.09%. Lastly, the addition of $\mathcal{L}_{\rm noise}$ results in further DSC enhancements by 0.44%, 0.29% and 2.18%. These results demonstrate the effectiveness of all components, particularly $\mathcal{L}_{\rm noise}$, which significantly enhances the performance across all three segmentation tasks ($p\leq 0.05$).
@@ 4.3.2 Ablation Studies of Temporal Ensembling
4.3.3 Hyper-parameter and Prototype-refined Strategy Analysis
We evaluate the influence of certain hyper-parameters and the efficacy of utilizing prototype-refined predictions on the ODOC task. The hyper-parameters of interest are the trade-off coefficients $\lambda_{3}$ and $\lambda_{4}$ in Eq. (17), which are employed to balance $\mathcal{L}_{\rm PRSA-main}$ and $\mathcal{L}_{\text{noise }}$. As shown in Fig. 4 Left, $\lambda_{3}$ is suitable to be set to 0.1 and 0.2, and $\lambda_{4}$ is suitable to be set to 0.01 and 0.05. The best performance is achieved when $\lambda_{3}=0.1$ and $\lambda_{4}=0.01$. We apply this optimal configuration to the other two tasks as well. The efficacy of utilizing prototype-refined predictions can be indirectly assessed by comparing “w/o Refinement” with “w/ Refinement” in Fig. 4 Right. With the increase in Iteration, the performance in both settings consistently improves and $\triangle D$ gradually increases from 3.67% to 5.70%. These results indicate that extra knowledge can be learned by the segmentation model through the application of prototype-refined predictions.
Figure 4: Left: Performance with varied trade-off coefficients $\lambda_{3}$ and $\lambda_{4}$. Right: Qualitative evaluation on the prototype-refined predictions. “w/ Refinement” and “w/o Refinement” respectively refer to the outputs of models employing ProCNS with and without using the prototype-refined strategy. $\Delta D$ denotes the DSC difference between “w/o Refinement” and “w/ Refinement”.
Prototype Variant DSCnormal-$\uparrow$ HD95normal-$\downarrow$
@@ 4.3.3 Hyper-parameter and Prototype-refined Strategy Analysis
Table 4: Comparison with state-of-the-art WSS methods on the FAZ, Polyp and ODOC segmentation tasks. $\text{UNet}_{pCE}$ represents the UNet model trained with sparse annotations, employing only the pCE loss. $\text{UNet}_{F}$ represents the model trained with full annotations, utilizing only the CE loss. “***” indicates $\rm p\leq 0.05$ in a Wilcoxon matched-pairs signed rank test when comparing ProCNS with the best-performing method in other WSS methods. The best results are in bold and the second-best results are underlined.
@@
DSCnormal-$\uparrow$ HD95normal-$\downarrow$ DSCnormal-$\uparrow$ HD95normal-$\downarrow$ DSCnormal-$\uparrow$ HD95normal-$\downarrow$ DSCnormal-$\uparrow$ HD95normal-$\downarrow$ DSCnormal-$\uparrow$ HD95normal-$\downarrow$
@@
Figure 5: Visualization of representative segmentation results from ProCNS and other SOTA WSS methods on the FAZ, Ployp and ODOC tasks.
4.3.4 Ablation Studies of Prototype
@@
Table 5: Validation experiments of ProCNS’s potential as a seamless integration plugin on the FAZ, Ployp and ODOC tasks. “***” represents $\rm p\leq 0.05$ in a Wilcoxon signed-rank test comparing the corresponding evaluation metric of the WSS method with and without ProCNS.
@@
DSCnormal-$\uparrow$ HD95normal-$\downarrow$ DSCnormal-$\uparrow$ HD95normal-$\downarrow$ DSCnormal-$\uparrow$ HD95normal-$\downarrow$
@@
TreeEnergy w/ ProCNS 93.22(+0.38) 7.36(-0.30) $*{}^{*}$ $*{}^{*}$ 87.93(+0.30) $*{}^{*}$
$\text{S}^{2}$ ME [41] 92.33 12.44 77.26 57.91 87.82 10.24
$\text{S}^{2}$ ME w/ ProCNS $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ 9.94(-0.30)
@@
4.4 Comparison with State-of-the-arts
To validate the superiority of the proposed algorithm, we compare ProCNS with a series of state-of-the-art WSS algorithms across the three distinct segmentation tasks. Among these compared methods, USTM [40] exploits consistency learning techniques to regularize the training process. S2L [31] and DMPLS [12] propagate semantic information from labeled pixels to unlabeled ones to generate pseudo-labels. GatedCRF [6] and TreeEnergy [4] employ low-dimensional features from the original images for additional supervision.
As shown in Table 4, ProCNS outperforms almost all other compared methods on all the three segmentation tasks, with the exception of ScribbleVC [42] in terms of DSC, which slightly outperforms ProCNS on OC segmentation. The superiority of ProCNS could be due to the collaborative synergy between the PRSA loss and ANPM. Specifically, the PRSA loss leverages the calibrated prototypes updated by ANPM’s noise masking mechanism to provide more precise guidance. On the other hand, ANPM utilizes the prediction refinement strategy of the PRSA loss to better perceive noisy regions. It is observed that the compared methods may be more suitable for segmenting relatively regular anatomical structures. When dealing with irregular pathological regions of interest, such as polyps, these methods exhibit subpar performance. In contrast, our approach exhibits superior generalization, not only excelling on the segmentation of the former but also delivering highly satisfactory performance on the latter. Specifically, for polyp segmentation, our proposed method significantly outperforms the second-best approach (EM), achieving a DSC improvement of 3.25% at $\rm p\leq 0.05$. On FAZ segmentation, compared over the second-best method (TreeEnergy), ProCNS achieves a noteworthy DSC improvement of 0.9% at $\rm p\leq 0.05$, with the difference from the fully-supervised DSC being merely 1.4%. Our method holds a 0.51% lead over the second-ranked method (EM) on OD segmentation at $\rm p\leq 0.05$. The average DSC of ODOC from ProCNS is 0.23% higher than that from the second-best method (ScribbleVC).
4.5 Verification of Seamless Plugin Integration
@@
In this subsection, we aim to experimentally validate our statement that ProCNS’s Main stage can serve as a seamlessly integrable plugin for other WSS methods, further enhancing model performance. Specifically, we replace the Initialization stage of ProCNS with four different WSS methods (including their models, losses, and training paradigms) and train them until convergence is achieved (obtaining preliminary segmentation models). Subsequently, we proceed with the Main stage. As tabulated in Table 5, it is evident that compared to the original methods, the approaches with ProCNS as a follow-up plugin consistently demonstrate varying degrees of performance improvements across the three tasks. Notably, the integration of ProCNS with pCE [14] results in a significant 3.68% DSC improvement on the FAZ task compared to the original method. Similarly, the combination of either TreeEnergy [4] or $\text{S}^{2}$ [41] with ProCNS respectively yields DSC improvements of 5.43% and 2.57% over the original counterparts. This confirms the capability of ProCNS to act as a follow-up plugin, aiding WSS methods by mitigating the impact of noise accumulation and overconfident prediction, thereby enhancing the final segmentation performance.

### v3

@@ 4.3.1 Ablation Studies of ProCNS
Table 2: Ablation study on the three key components in ProCNS regarding DSC. Here, $\checkmark$ indicates the corresponding component is applied. “*” represents $\rm p\leq 0.05$ in a Wilcoxon matched-pairs signed rank test comparing the DSC of the pertinent component before and after application. The column for ODOC reports the average DSC of OD and OC. The best results are in bold.
@@
$\checkmark$ 92.38 87.91 79.46
$\checkmark$ $\checkmark$ 93.30 88.05 80.55
$\checkmark$ $\checkmark$ $\checkmark$ 93.74 88.32 82.73
Table 3: Ablation study on the initial PRSA loss regarding DSC. The column for ODOC reports the average DSC of OD and OC.
Method FAZ ODOC Polyp
[…]
ProCNS 93.74 88.32 82.21
To ascertain the efficacy of the three key components $\mathcal{L}_{\rm PRSA}$, ANPM and $\mathcal{L}_{\rm noise}$ within ProCNS, we perform a series of ablation studies. In this context, the baseline is defined as the segmentation model that is trained with sparse annotations and solely employs the pCE loss.
Then, we incrementally incorporate the three components into the baseline. Table 2 presents the quantitative results from the ablation analyses on the three tasks in terms of DSC. Compared to the baseline, $\mathcal{L}_{\rm PRSA}$ respectively improves model performance by 2.36%, 0.52% and 0.47% for the FAZ, ODOC and polyp segmentation tasks. The integration of ANPM yields DSC improvements of 0.92%, 0.14% and 1.09%. Lastly, the addition of $\mathcal{L}_{\rm noise}$ results in further DSC enhancements by 0.44%, 0.29% and 2.18%. These results demonstrate the effectiveness of all components, particularly $\mathcal{L}_{\rm noise}$, which significantly enhances the performance across all three segmentation tasks ($p\leq 0.05$).
4.3.2 Ablation Studies of Initial PRSA Loss
To ascertain the efficacy of utilizing the initial PRSA loss at the Initialization stage, we report the results on the FAZ, ODOC and polyp tasks with and without the initial PRSA loss, as shown in Table 3. All results with the initial PRSA loss outperform those without it, indicating that the initial PRSA loss is beneficial for model performance. This may be attributed to the effective utilization of the high-level semantic correlation between unlabeled and sparsely-annotated regions, which aids the preliminary model in capturing boundary features and achieves a more reliable preliminary model for the Main stage, consequently assisting the final model in more accurately segmenting targets.
[…]
Figure 4: Performance with varied trade-off coefficients $\lambda_{3}$ and $\lambda_{4}$.
@@ 4.3.3 Ablation Studies of Temporal Ensembling
4.3.4 Hyper-parameter and Prototype-refined Strategy Analysis
Figure 5: Qualitative evaluation on the prototype-refined predictions. “w/ Refinement” and “w/o Refinement” respectively refer to the outputs of models employing ProCNS with and without using the prototype-refined strategy. $\Delta D$ denotes the DSC difference between “w/o Refinement” and “w/ Refinement”.
We evaluate the influence of certain hyper-parameters and the efficacy of utilizing prototype-refined predictions on the ODOC task. The hyper-parameters of interest are the trade-off coefficients $\lambda_{3}$ and $\lambda_{4}$ in Eq. (17), which are employed to balance $\mathcal{L}_{\rm PRSA-main}$ and $\mathcal{L}_{\text{noise }}$. As shown in Fig. 4, $\lambda_{3}$ is suitable to be set to 0.1 and 0.2 and $\lambda_{4}$ is suitable to be set to 0.01 and 0.05. The best performance is achieved when $\lambda_{3}=0.1$ and $\lambda_{4}=0.01$. We apply this optimal configuration to the other two tasks as well. The efficacy of utilizing prototype-refined predictions can be indirectly assessed by comparing “w/o Refinement” with “w/ Refinement” in Fig. 5. With the increase in Iteration, the performance in both settings consistently improves and $\triangle D$ gradually increases from 3.67% to 5.70%. These results indicate that extra knowledge can be learned by the segmentation model through the application of prototype-refined predictions.
@@ 4.3.4 Hyper-parameter and Prototype-refined Strategy Analysis
4.3.5 Ablation Studies of Prototype
Table 5: Comparison with state-of-the-art WSS methods on computational burden.
Method
Training Time
(s/epoch)
Testing Effiency
(samples/s)
$\text{UNet}_{pCE}$ 4.71 39
EM [54] 4.67 39
Random Walks [14] 2.62 38
AC [55] 2.71 39
GatedCRF [6] 4.86 35
S2L [38] 2.98 38
SeL [11] 6.24 38
USTM [56] 16.34 38
DMPLS [13] 4.22 33
TreeEnergy [4] 10.80 36
$\text{S}^{2}$ [57] 7.72 38
ScribbleVC [58] 28.62 12
ProCNS (Ours) 10.75 38
Table 6: Comparison with state-of-the-art WSS methods on the FAZ, Polyp and ODOC segmentation tasks. $\text{UNet}_{pCE}$ represents the UNet model trained with sparse annotations, employing only the pCE loss. $\text{UNet}_{F}$ represents the model trained with full annotations, utilizing only the CE loss. “*” indicates $\rm p\leq 0.05$ in a Wilcoxon matched-pairs signed rank test when comparing ProCNS with the best-performing method in other WSS methods. The best results are in bold and the second-best results are underlined.
@@ 4.3.5 Ablation Studies of Prototype
[…]
@@
Table 7: Comparison with state-of-the-art WSS methods on the WT, Nuclei and CM segmentation tasks. $\text{UNet}_{pCE}$ represents the UNet model trained with sparse annotations, employing only the pCE loss. $\text{UNet}_{F}$ represents the model trained with full annotations, utilizing only the CE loss. “*” indicates $\rm p\leq 0.05$ in a Wilcoxon matched-pairs signed rank test when comparing ProCNS with the best-performing method in other WSS methods. The best results are in bold and the second-best results are underlined.
Method BraTS2019 WO ACDC
WT Nuclei RV MYO LV Avg.
[…]
$\text{UNet}_{pCE}$ 72.94$\pm$25.43 14.10$\pm$15.35 72.63$\pm$5.42 90.28$\pm$42.37 64.13$\pm$26.25 131.38$\pm$41.75 65.72$\pm$17.51 120.80$\pm$36.13 79.04$\pm$20.99 85.50$\pm$58.87 69.63 112.56
EM [54] 73.26$\pm$25.52 17.02$\pm$20.95 72.30$\pm$8.43 70.25$\pm$40.43 81.72$\pm$26.05 14.65$\pm$32.02 80.07$\pm$14.95 16.76$\pm$31.30 87.85$\pm$17.93 13.45$\pm$30.53 83.22 14.95
Random Walks [14] 78.05$\pm$23.71 19.44$\pm$25.18 69.36$\pm$6.56 75.03$\pm$38.79 77.50$\pm$27.48 7.87$\pm$9.77 66.64$\pm$22.88 6.81$\pm$7.47 78.80$\pm$30.93 3.58$\pm$6.98 74.31 6.09
AC[55] 74.33$\pm$23.68 16.19$\pm$17.52 73.14$\pm$5.06 91.45$\pm$46.97 71.38$\pm$27.11 102.47$\pm$52.42 68.61$\pm$16.87 114.92$\pm$43.36 77.87$\pm$23.39 63.08$\pm$59.78 72.62 93.49
GatedCRF[6] 77.15$\pm$23.45 17.81$\pm$22.91 75.34$\pm$4.85 76.08$\pm$44.35 82.05$\pm$25.48 7.25$\pm$12.60 80.82$\pm$15.01 4.84$\pm$10.59 87.65$\pm$17.71 4.54$\pm$9.77 83.51 5.54
S2L [38] 70.35$\pm$23.53 17.48$\pm$20.53 74.08$\pm$5.43 73.56$\pm$39.23 77.67$\pm$25.31 59.22$\pm$52.45 70.92$\pm$16.72 73.55$\pm$50.93 78.39$\pm$21.76 65.89$\pm$53.53 75.66 66.22
SeL [11] 76.59$\pm$25.92 14.64$\pm$19.95 74.53$\pm$5.64 71.02$\pm$39.51 78.84$\pm$29.38 6.24$\pm$9.42 75.76$\pm$16.50 56.37$\pm$58.93 85.33$\pm$20.34 3.96$\pm$6.73 79.98 22.19
USTM [56] 71.43$\pm$21.09 15.74$\pm$15.56 72.52$\pm$6.12 75.26$\pm$42.46 78.52$\pm$26.13 56.75$\pm$58.16 77.54$\pm$14.74 50.15$\pm$52.63 84.37$\pm$19.05 45.59$\pm$56.97 80.15 50.83
DMPLS [13] 73.24$\pm$25.11 14.83$\pm$17.71 73.26$\pm$6.53 80.48$\pm$41.65 81.62$\pm$25.52 5.83$\pm$9.39 82.49$\pm$13.24 6.91$\pm$15.27 88.67$\pm$17.41 4.04$\pm$8.92 84.26 5.59
TreeEnergy[4] 77.90$\pm$22.56 14.51$\pm$16.25 71.69$\pm$6.38 90.27$\pm$48.94 82.06$\pm$24.39 12.42$\pm$24.39 80.11$\pm$13.55 22.33$\pm$33.15 87.45$\pm$16.77 16.01$\pm$30.88 83.20 16.92
$\text{S}^{2}$ [57] 75.72$\pm$25.09 14.78$\pm$19.45 73.14$\pm$5.06 91.45$\pm$46.97 83.10$\pm$24.17 6.09$\pm$8.95 76.51$\pm$15.96 53.78$\pm$56.26 87.83$\pm$16.66 3.69$\pm$6.97 82.48 21.19
ScribbleVC [58] 78.33$\pm$25.63 13.68$\pm$19.09 73.54$\pm$5.95 71.72$\pm$41.85 82.84$\pm$25.35 5.49$\pm$7.66 80.28$\pm$12.77 6.35$\pm$12.51 88.56$\pm$16.37 5.65$\pm$14.21 83.89 5.83
ProCNS (Ours) 79.49$\pm$23.33 14.34$\pm$19.96 76.11$\pm$4.86 69.11$\pm$40.15 83.83$\pm$24.78 7.52$\pm$15.27 81.66$\pm$14.30 5.51$\pm$10.78 88.70$\pm$17.05 3.97$\pm$9.15 84.73 5.66
$\text{UNet}_{F}$ 82.25$\pm$20.13 15.34$\pm$23.35 79.30$\pm$5.57 63.86$\pm$39.74 84.88$\pm$25.30 4.51$\pm$7.65 83.76$\pm$15.22 4.07$\pm$8.27 89.40$\pm$18.25 3.81$\pm$8.74 86.01 4.13
@@
Figure 6: Visualization of representative segmentation results from ProCNS and other SOTA WSS methods on the FAZ, Ployp, ODOC, WT, Nuclei and CM tasks.
4.4 Comparison with State-of-the-arts
To validate the superiority of the proposed algorithm, we compare ProCNS with a series of state-of-the-art WSS algorithms across the six distinct segmentation tasks. Among these compared methods, USTM [56] exploits consistency learning techniques to regularize the training process. Seminar Learning (SeL) [11] distills fine-grained semantic infromation from labeled pixels to guide the training of the final model. S2L [38] and DMPLS [13] propagate semantic information from labeled pixels to unlabeled ones to generate pseudo-labels. GatedCRF [6] and TreeEnergy [4] employ low-dimensional features from the original images for additional supervision.
In terms of computational burden, in Table 5 we present the training time and testing efficiency for various compared methods on the FAZ task. It is important to note that, for a fairness purpose, these metrics are provided under identical hardware and communication conditions. The results indicate that ProCNS has a slightly increased training time compared to certain WSS methods while maintaining comparable testing efficiency.
In terms of model performance, as shown in Table 6 and Table 7, ProCNS outperforms almost all other compared methods on all the six segmentation tasks. The superiority of ProCNS could be due to the collaborative synergy between the PRSA loss and ANPM. Specifically, the PRSA loss leverages the calibrated prototypes updated by ANPM’s noise masking mechanism to provide more precise guidance. On the other hand, ANPM utilizes the prediction refinement strategy of the PRSA loss to better perceive noisy regions. It is observed that the compared methods may be more suitable for segmenting relatively regular anatomical structures. When dealing with irregular pathological regions of interest, such as polyps, these methods exhibit subpar performance. In contrast, our approach exhibits superior generalization, not only excelling on the segmentation of the former but also delivering highly satisfactory performance on the latter. Specifically, for polyp segmentation, our proposed method significantly outperforms the second-best approach (EM), achieving a DSC improvement of 3.25% at $\rm p\leq 0.05$. On FAZ segmentation, compared over the second-best method (TreeEnergy), ProCNS achieves a noteworthy DSC improvement of 0.9% at $\rm p\leq 0.05$, with the difference from the fully-supervised DSC being merely 1.4%. Our method holds a 0.51% lead over the second-ranked method (EM) on OD segmentation at $\rm p\leq 0.05$. The average DSC of ODOC from ProCNS is 0.23% higher than that from the second-best method (ScribbleVC). For WT segmentation, our method holds a 1.16% lead over the second-ranked method (ScribbleVC) at $\rm p\leq 0.05$. For nuclei segmentation, our method significantly outperforms the second-best approach (GatedCRF), achieving a DSC improvement of 0.77% at $\rm p\leq 0.05$. On the CM segmentation task, our method holds a 0.73% lead over the second-ranked method ($\text{S}^{2}$) for the RV structure at $\rm p\leq 0.05$. The average DSC of RV, MYO and LV from ProCNS is 0.47% higher than that from the second-best method (DMPLS).
4.5 Verification of Seamless Plugin Integration
@@
Table 8: Validation experiments of ProCNS’s potential as a seamless integration plugin on the FAZ, Ployp and ODOC tasks. “*” represents $\rm p\leq 0.05$ in a Wilcoxon signed-rank test comparing the corresponding evaluation metric of the WSS method with and without ProCNS.
@@
[…]
@@
TreeEnergy w/ ProCNS 93.22(+0.38) 7.36(-0.30) 82.37(+5.43) 51.45(-0.56) 87.93(+0.30) 10.19(-0.87)
$\text{S}^{2}$ [57] 92.33 12.44 77.26 57.91 87.82 10.24
$\text{S}^{2}$ w/ ProCNS 92.93(+0.60) 6.89(-5.55) 81.41(+4.15) 54.01(-3.90) 88.33(+0.51) 9.94(-0.30)
@@
Figure 7: Visualization of pseudo-labels generated by SAM-Med2D [60]. The overlaid green and red points respectively denote target and background prompts.
@@
In this subsection, we aim to experimentally validate our statement that ProCNS’s Main stage can serve as a seamlessly integrable plugin for other WSS methods, further enhancing model performance. Specifically, we replace the Initialization stage of ProCNS with four different WSS methods (including their models, losses and training paradigms) and train them until convergence is achieved (obtaining preliminary segmentation models). Subsequently, we proceed with the Main stage. As tabulated in Table 8, it is evident that compared to the original methods, the approaches with ProCNS as a follow-up plugin consistently demonstrate varying degrees of performance improvements across the three tasks. Notably, the integration of ProCNS with pCE [15] results in a significant 3.68% DSC improvement on the FAZ task compared to the original method. Similarly, the combination of either TreeEnergy [4] or $\text{S}^{2}$ [57] with ProCNS respectively yields DSC improvements of 5.43% and 2.57% over the original counterparts. This confirms the capability of ProCNS to act as a follow-up plugin, aiding WSS methods by mitigating the impact of noise accumulation and overconfident prediction, thereby enhancing the final segmentation performance.
Table 9: Performance comparisons (DSC) with and without utilizing pseudo-labels generated by SAM-Med2D on the FAZ and polyp tasks.
Method FAZ Polyp
SAM-Med2D w/ ProCNS 93.30 79.04
ProCNS 93.74 82.21
4.6 Exploratory Analysis of Incorporating Foundation Models
Benefiting from the development of SAM [61], several studies [62] have emerged that utilize foundation models to assist WSS. We assess the impact of integrating foundation models into ProCNS. Specifically, we exclude the Initialization stage and utilize predictions from a foundation model as the initial pseudo-labels in the Main stage. Recent research [60, 63, 64] indicates that directly applying the pretrained SAM to medical image segmentation is suboptimal due to the significant domain gap between natural and medical images. Therefore, we employ SAM’s medical variant, e.g., SAM-Med2D [60], to generate pseudo-labels for the FAZ and polyp tasks (as depicted in Fig. 7). As shown in Table 6 and Table 9, for those two FAZ and polyp tasks, utilizing pseudo-labels generated by SAM-Med2D (SAM-Med2D w/ ProCNS) yields better performance than most WSS methods. Despite its performance still lagging behind the original design of ProCNS, the results still indicate that leveraging foundation models to assist WSS is a promising direction.
4.7 Analysis on the Impact of Annotation Sparsity